# Axi-majoron for almost everything

### v2

@@ 2 The Model
$S_{\rm M}$ $=$ $\int d^{4}x\sqrt{-{\tilde{g}}}\left[\frac{1}{2}{\tilde{g}}^{\mu\nu}\left(\partial_{\mu}\chi\partial_{\nu}\chi+\partial_{\mu}h\partial_{\nu}h+\sum_{a}\partial_{\mu}\phi^{a}\partial_{\nu}\phi^{a}\right)-{\tilde{V}}\right]\ ,$ (2.8)
where in Eq. (2.7) we redefined $\Phi$ as
@@
$\left(\xi_{+}-i\xi_{-}\right)e^{i2\theta_{0}}=\sqrt{\xi_{+}^{2}+\xi_{-}^{2}}\equiv\xi_{a}$ (2.10)
and $\phi^{a}=\left(\phi_{r},\phi_{i}\right)^{T}$ in Eq. (2.8). We define a frame-function $\Omega$ as
$\Omega$$\equiv$ $\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{h}h^{2}}{M_{\rm P}^{2}}+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\left(1+\alpha\cos(2\theta)\right)$ (2.11)
$=$ $\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{h}h^{2}}{M_{\rm P}^{2}}+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P}^{2}}\ ,$
where $\alpha\equiv\xi_{a}/\xi_{\phi}$, $\xi_{r}\equiv\xi_{\phi}+\xi_{a}$, $\xi_{i}\equiv\xi_{\phi}-\xi_{a}$ ($\xi_{r}\geq\xi_{i}$). We assume
@@
in order for $\Omega$ to be positive definite even in the large field region, $\xi_{\phi}^{1/2}\phi>M_{\rm P}$, which will be of our interest for a primordial inflation. Without loss of generality, the scalar potential ${\tilde{V}}_{S}$ can be expressed as  [^7] The presence of the second Higgs doublet makes the shape of the potential complicated, but such details are not important in the discussion of the present work.
${\tilde{V}}_{S}(\chi,h,\phi)$ $=$ $\frac{\lambda_{\chi}^{\prime}}{4}\chi^{4}+\frac{\lambda_{\phi}^{\prime}}{4}\left(\phi^{2}-\zeta_{\phi\chi}\chi^{2}\right)^{2}+\frac{\lambda_{h}}{4}\left(h^{2}+\zeta_{h\phi}\phi^{2}-\zeta_{h\chi}\chi^{2}\right)^{2}\ .$ (2.13)
Hence, once the vacuum expectation value (VEV) of $\chi$ is fixed so that $\xi_{\chi}^{1/2}\langle\chi\rangle=M_{\rm P}\simeq 2.44\times 10^{18}{\rm GeV}$(see Ref. [9] for example for dynamical generation of Planck scale in a scale invariant theory), the VEVs of $\phi$ and $h$ at low energy can be fixed to
$\phi_{0}$ $=$ $\sqrt{\zeta_{\phi\chi}\xi_{\chi}^{-1}}M_{\rm P}=\mathcal{O}(10^{9-12})\ {\rm GeV}\ ,$ (2.14)
$h_{0}$ $=$ $\sqrt{\zeta_{h\chi}\xi_{\chi}^{-1}M_{\rm P}^{2}-\zeta_{h\phi}\phi_{0}^{2}}\simeq 246\ {\rm GeV}\ ,$ (2.15)
and ${\tilde{V}}_{S}$ has the familiar form,
${\tilde{V}}_{S}(h,\phi)=V_{0}+\frac{\lambda_{h}}{4}\left(h^{2}-h_{0}^{2}\right)^{2}+\frac{\lambda_{\phi}^{\prime}}{4}\left(\phi^{2}-\phi_{0}^{2}\right)^{2}\ .$ (2.16)
Note that, if $\lambda_{h\phi}=0$(or equivalently $\zeta_{h\phi}=0$) which we assume in this work [^8] Actually, it is enough to have $\lambda_{h\phi}\ll\left(\lambda_{h\chi}/\xi_{\chi}\right)\times\left(M_{\rm P}/\phi_{0}\right)^{2}$ which is not a special condition., the VEVs are determined by the couplings to $\chi$ only, i.e., $\chi$ is responsible for all scales.
In the theory characterized by the action in Eqs. (2.1) and (2.2), Planck scale is determined by the sum of the VEVs of all scalar fields. The sum, however, should be dominated by the VEV of $\chi$, since the other fields are supposed to have much smaller VEVs at low energy. The determination of Planck scale implies breaking of scale-invariance. Since there are no explicit symmetry-breaking terms, this should be achieved by spontaneous symmetry-breaking. In Ref. [9], it was shown that Planck scale can emerge naturally in scale-invariant theories even without CW-type dimensional transmutation. This is due to the fact that the conserved scale-current has a kernel which eventually settles down to a constant after some general expansion in a very early epoch, at least well before the last about $60$ inflationary $e$-folds. Scale-anomaly can compromise this picture, but it was shown that, once the renormalization scale, which usually appears as a dimensionful parameter, is replaced by a field, scale symmetry becomes exact, and the existence of a constant kernel remains valid [36]. The constant value of the kernel defines an ellipse in the multi-dimensional scalar-field space 999 Quantum effects on $\xi_{i}$ can change the elliptic configuration [9], but the impact on our discussion is minor. Hence, we ignore the effects in this work., and eventually determines the Planck scale at low energy, once all fields settle down to their true vacuum position.
Applying the analysis of Ref. [9] to our scenario, when $h\lll\chi,\phi_{r},\phi_{i}$ in the very early epoch including inflation era, we are left with a three-dimensional field space. In terms of new field variables defined as  [^10] The index $\alpha$ should not be confused with $\alpha$ defined in Eq. (2.11).
@@
with $\alpha=(1,2,3)$, $\xi_{\alpha}=\left(\xi_{\chi},\xi_{r},\xi_{i}\right)$, and $\varphi_{\alpha}=\left(\chi,\phi_{r},\phi_{i}\right)$, the constant kernel is given by
$K=\frac{1}{2}\sum_{\alpha}\kappa_{\alpha}\varphi_{\alpha}^{2}\ ,$ (2.18)
where $\kappa_{\alpha}\equiv 1+6\xi_{\alpha}$. If the initial condition and the dynamics of fields are such that
$\kappa_{\chi}\chi^{2}\gg\kappa_{r}\phi_{r}^{2}+\kappa_{i}\phi_{i}^{2}$ (2.19)
during the whole evolution of fields, $K$ would always be dominated by the contribution of $\chi$. In terms of $X_{\alpha}$, the condition Eq. (2.19) corresponds to
@@
For $\xi_{a}\ll\xi_{\phi}$(i.e., $\alpha\ll 1$) which will be of our interest, $\xi_{2}\approx\xi_{3}$ and we can take $\xi_{\alpha}$s such as
@@
as the initial condition for the primordial inflation responsible for our universe. If ${\tilde{V}}_{S}$ is dominated by the $\lambda_{\phi}\phi^{4}$-term during inflation, in the slow-roll limit of the primordial inflation, the evolution equation in the Jordan frame is approximated as  [^11] Actually, since the dynamics is constrained to be on the ellipse defined in Eq. (2.18), one of the equations in Eq. (2.26) is redundant. [9]
@@
Hence, one can obtain sufficiently large amount of $e$-folds if $X_{2}(0)\gg X_{1}(0)$. Note that, since $X_{1}$ barely changes for a sufficiently small $\xi_{1}$ while $X_{2,3}$ becomes negligible in the end, it is possible to assume an approximation of a fixed Planck scale whose value given by
@@
Then, the physics becomes the same as that of non-minimal gravitational couplings of scalar fields other than $\chi$ with the fixed Planck scale. We take this approximation in the subsequent discussion to describe the dynamics $X_{2,3}$ in the more transparent and clear, conventional way.
Once the Planck scale is (nearly) fixed in this way, we can follow the conventional way of converting the Jordan frame to the Einstein frame. The frame-function in Eq. (2.11) with the Higgs field ignored is now expressed as
$\Omega\equiv 1+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\left(1+\alpha\cos(2\theta)\right)=1+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P}^{2}}\ .$ (2.35)
@@
while the action of matter sector is given by
$S_{\rm M}=\int d^{4}x\sqrt{-g}\left\{\frac{1}{2\Omega}\left[\left(\partial h\right)^{2}+\sum_{a}\left(\partial\phi^{a}\right)^{2}\right]-\frac{{\tilde{V}}}{\Omega^{2}}\right\}\ .$ (2.38)
The action for the scalar fields part can now be written as
@@
where the symmetric field-space metric $\mathcal{K}_{ab}$ is given by
$\mathcal{K}_{ab}=\Omega^{-2}\left[\Omega\delta_{ab}+\frac{3M_{\rm P}^{2}}{2}\frac{\partial\Omega}{\partial\phi^{a}}\frac{\partial\Omega}{\partial\phi^{b}}\right]\equiv\Omega^{-2}K_{ab}$ (2.40)
and $U\equiv{\tilde{V}}_{S}(h,\phi)/\Omega^{2}$. At low energy with $h,\phi\lll M_{\rm P}$, $\Omega\to 1$ and the kinetic terms return to the canonical form. The shape of the potential $U$ is shown in Fig. 1.
@@
The non-minimal symmetry-breaking gravitational interaction, $\xi_{a}$-term contributes to the mass of the axi-majoron, $a_{\phi}\equiv\theta\phi$, the angular mode of $\Phi$ field for a given non-zero $\phi$. The low-energy definition of axi-majoron differs due to the symmetry-breaking non-zero VEVs of the PQ-charged two Higgses, but such a change does not affect our argument. In the Einstein frame, in the vicinity of true vacuum position, the mass-squared can be read as
$m_{a}^{2}\equiv\frac{\partial^{2}U}{\phi^{2}\partial\theta^{2}}=8\alpha\xi_{\phi}\frac{U}{M_{\rm P}^{2}}\ .$ (2.41)
Hence, depending on $\alpha\ \xi_{\phi}$, it can be comparable to or smaller than the expansion rate. Note that, even though it is from a symmetry-breaking term, this contribution to the axi-majoron mass does not disturb the axion solution, since it becomes negligible at low energy, but helps suppress the iso-curvature perturbations and play the key role for baryogenesis after inflation, as discussed in the subsequent sections.

### v3

@@ 2 The Model
$\tilde{\phi}_{0}^{2}\simeq\zeta_{\phi\chi}+\frac{\lambda_{\phi h}}{2\lambda_{\phi}^{\prime}\sqrt{\zeta_{1}\zeta_{2}}}\left(\zeta_{h\phi}+\frac{\lambda_{\phi h}}{8\lambda_{h}\sqrt{\zeta_{1}\zeta_{2}}}\right)\simeq\zeta_{\phi\chi}\ ,$ (2.14)
and the tree-level cosmological constant(CC) term is given by
$U_{0}$ $\simeq$ $\frac{M_{\rm P}^{4}}{4\xi_{\chi}^{2}}\left\{\lambda_{\chi}^{\prime}+\lambda_{\phi}^{\prime}\zeta_{\phi\chi}^{2}\left[1-\frac{1+\frac{\lambda_{\phi h}\zeta_{h\chi}}{4\lambda_{\phi}^{\prime}\sqrt{\zeta_{1}\zeta_{2}}\zeta_{\phi\chi}}}{1-\frac{\lambda_{\phi h}}{2\lambda_{\phi}^{\prime}\sqrt{\zeta_{1}\zeta_{2}}}\left(\zeta_{h\phi}+\frac{\lambda_{\phi h}}{8\lambda_{h}\sqrt{\zeta_{1}\zeta_{2}}}\right)}\right]\right\}$ (2.15)
$\simeq$ $\frac{M_{\rm P}^{4}}{4\xi_{\chi}^{2}}\left[\lambda_{\chi}^{\prime}-\frac{\lambda_{\phi h}}{2\lambda_{\phi}^{\prime}\sqrt{\zeta_{1}\zeta_{2}}}\left(\zeta_{h\phi}+\frac{\lambda_{\phi h}}{8\lambda_{h}\sqrt{\zeta_{1}\zeta_{2}}}\right)\lambda_{\phi}^{\prime}\zeta_{\phi\chi}^{2}\right]$
There would be quantum corrections to the CC-term, but we will not discuss them, since it can be canceled by a proper choice of $\lambda_{\chi}^{\prime}$ and we are not trying to solve the CC-problem in this work.
We may express $U$ at low energy in a form,
$U(h,\phi)$ $=$ $U_{0}+\frac{\lambda_{\phi}^{\prime}}{4}\left(\phi^{2}-\phi_{0,\chi}^{2}\right)^{2}+\frac{\lambda_{h}}{4}\left(\zeta_{1}h_{1}^{2}+\zeta_{2}h_{2}^{2}-\zeta_{h\phi}\phi^{2}-h_{0,\chi}^{2}\right)^{2}$ (2.16)
-- $\frac{1}{4}\lambda_{\phi h}\phi^{2}h_{1}h_{2}\cos\left(2\theta-\theta_{1}+\theta_{2}\right)\ ,$
where
$\phi_{0,\chi}$$\equiv$ $\sqrt{\zeta_{\phi\chi}\xi_{\chi}^{-1}}M_{\rm P}\ ,$ (2.17)
$h_{0,\chi}$$\equiv$ $\sqrt{\zeta_{h\chi}\xi_{\chi}^{-1}}M_{\rm P}\ .$ (2.18)
Also, Eq. (2.13) with $\tilde{\phi}$ replaced to $\tilde{\phi}_{0}$ gives the VEV of $h_{i}$ as
$h_{i,0}^{2}=\frac{M_{\rm P}^{2}}{2\zeta_{i}\xi_{\chi}}\left[\zeta_{h\chi}+\left(\zeta_{h\phi}+\frac{\lambda_{\phi h}}{4\lambda_{h}\sqrt{\zeta_{1}\zeta_{2}}}\right)\tilde{\phi}_{0}^{2}\right]\lesssim\mathcal{O}(10^{2}){\rm GeV}$ (2.19)
which should satisfy $v_{h}\equiv\sqrt{h_{1,0}^{2}+h_{2,0}^{2}}\simeq 246{\rm GeV}$ for the correct electroweak physics. This implies that the quartic couplings between $\Phi$ and $H_{i}$ are constrained as $\lambda_{\phi H_{i}}\lesssim\left(h_{i,0}/\phi_{0}\right)^{2}$. The radiative corrections to the coupling between $\Phi$ and $H_{2}$, caused by the couplings $y_{N}$ and $y_{\nu}$, are also constrained so as to satisfy
$\mathcal{O}(10^{-2})\times\left(y_{\nu}\right)_{\alpha i}^{2}y_{N_{i}}^{2}\lesssim\left(\frac{v_{h}}{\phi_{0}}\right)^{2}\sin^{2}\beta$ (2.20)
where the subscripts ‘αi’ are flavor indices of left-handed and right-handed neutrino species, respectively, $\tan\beta\equiv h_{2,0}/h_{1,0}$, and the mass-matrix of right-handed neutrinos was assumed to be diagonal. Combined with the seesaw formula in Eq. (3.64), the condition constrains $y_{N_{i}}$ as
$y_{N_{i}}\lesssim\mathcal{O}(10^{-5})\times\left(\frac{0.05{\rm eV}}{m_{\nu_{\alpha}}}\right)^{1/3}\left(\frac{10^{12}{\rm GeV}}{\phi_{0}}\right)\sin^{\frac{4}{3}}\beta$ (2.21)
which is relevant for right-handed neutrino flavors associated with the largest $\left(y_{\nu}\right)_{\alpha i}$. It can be satisfied either trivially or marginally, depending on other parameters, as will become clear in the subsequent discussion. Also, if the contributions of $\phi_{0}$ and $\chi_{0}$ to $h_{i,0}$ (see Eq. (2.19)) have different signs, the constraint can become weaker. Hence, we ignore it in the discussion from now on.
In this work, we are interested in the case of $|\Phi|\gtrsim M_{\rm P}$ while $|H_{i}|\lll M_{\rm P}$ initially some time well before the last about $60$ $e$-folds of the primordial inflation responsible for our Universe777 We will discuss else where the case where initially $H_{1}$ and $H_{2}$ have very large field values during inflation. . In such a circumstance, the two Higgs doublets do not play a critical role in our discussion [^8] For $\langle|H_{1,2}|\rangle\ll M_{\rm P}$, the nonminimal gravitational couplings of the two Higgs-doublets can be essentially turned off(see Eq. (2.7)). Then, the electroweak symmetry-breaking is expected to take place as the case in the standard cosmology as long as the final reheating temperature after inflation is higher than the electroweak scale. . They are relevant only for low energy phenomenology which is separable from our discussion in this work. Also, the details of the vacuum structure of the electroweak sector does not affect our argument. Hence, we omit Higgs fields in the subsequent discussion except in some relevant places.
In regard of relevant scalar fields and their interactions only, the simplified model is given by the following action:
$S_{\rm G}$ $=$ $-\frac{1}{2}\int d^{4}x\sqrt{-{\tilde{g}}}{\tilde{R}}\left[\xi_{\chi}\chi^{2}+2\xi_{\phi}|\Phi|^{2}+\xi_{+}\left(\Phi^{2}+{\rm c.c.}\right)-i\xi_{-}\left(\Phi^{2}-{\rm c.c.}\right)\right]\ ,$ (2.22)
$S_{\rm M}$ $=$ $\int d^{4}x\sqrt{-{\tilde{g}}}\left[{\tilde{g}}^{\mu\nu}\left(\frac{1}{2}\partial_{\mu}\chi\partial_{\nu}\chi+\partial_{\mu}\Phi^{\dagger}\partial_{\nu}\Phi\right)+\mathcal{L}_{\rm ss}-{\tilde{V}}\right]\ ,$ (2.23)
where
${\tilde{V}}=\frac{\lambda_{\chi}}{4}\chi^{4}+\lambda_{\phi}|\Phi|^{4}-\frac{1}{2}\lambda_{\chi\phi}\chi^{2}|\Phi|^{2}\ .$ (2.24)
Redefining $\Phi$ as
@@
$\left(\xi_{+}-i\xi_{-}\right)e^{i2\theta_{0}}=\sqrt{\xi_{+}^{2}+\xi_{-}^{2}}\equiv\xi_{a}\ ,$ (2.26)
we can express the action as
$S_{\rm G}$ $=$ $-\frac{1}{2}\int d^{4}x\sqrt{-{\tilde{g}}}{\tilde{R}}M_{\rm P}^{2}\Omega\ ,$ (2.27)
$S_{\rm M}$ $=$ $\int d^{4}x\sqrt{-{\tilde{g}}}\left[\frac{1}{2}{\tilde{g}}^{\mu\nu}\left(\partial_{\mu}\chi\partial_{\nu}\chi+\sum_{a}\partial_{\mu}\phi^{a}\partial_{\nu}\phi^{a}\right)+\mathcal{L}_{\rm ss}-{\tilde{V}}\right]\ ,$ (2.28)
where in Eq. (2.27) the frame-function $\Omega$ is now given by
$\Omega\equiv\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\left(1+\alpha\cos(2\theta)\right)=\frac{\xi_{\chi}\chi^{2}}{M_{\rm P}^{2}}+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P}^{2}}$ (2.29)
with $\alpha\equiv\xi_{a}/\xi_{\phi}$, $\xi_{r}\equiv\xi_{\phi}+\xi_{a}$, $\xi_{i}\equiv\xi_{\phi}-\xi_{a}$, and $\phi^{a}\equiv\left(\phi_{r},\phi_{i}\right)^{T}$ in Eq. (2.28). We assume
@@
in order for $\Omega$ to be positive definite even in the large field region, $\xi_{\phi}^{1/2}\phi>M_{\rm P}$, which will be of our interest for a primordial inflation.
Applying the analysis of Ref. [8] to our scenario, when $h_{i}\lll\chi,\phi_{r},\phi_{i}$ in the very early epoch including inflation era, we are left with a three-dimensional field space $\varphi_{\alpha}=\left(\chi,\phi_{r},\phi_{i}\right)$. The constant kernel is given by
$K=\frac{1}{2}\sum_{\alpha}\kappa_{\alpha}\varphi_{\alpha}^{2}\ ,$ (2.31)
where $\kappa_{\alpha}\equiv 1+6\xi_{\alpha}$ with $\xi_{\alpha}=\left(\xi_{1},\xi_{2},\xi_{3}\right)=\left(\xi_{\chi},\xi_{r},\xi_{i}\right)$. It is of the same form as the case when $\xi_{\alpha}$s are pure numerical constants. If the initial condition and the dynamics of fields are such that
$\kappa_{\chi}\chi^{2}\gg\kappa_{r}\phi_{r}^{2}+\kappa_{i}\phi_{i}^{2}$ (2.32)
during the whole evolution of fields, $K$ would always be dominated by the contribution of $\chi$. In terms of new field variables defined as
@@
the condition Eq. (2.32) corresponds to
@@
For $\xi_{a}\ll\xi_{\phi}$ which will be of our interest, $\xi_{2}\approx\xi_{3}$ and we can take $\xi_{\alpha}$s such as
@@
as the initial condition for the primordial inflation responsible for our Universe. Also, as long as $c_{\phi}\chi\ll\phi$ is satisfied at least during inflation, one can treat $\xi_{\phi}$ as a constant. We assume this is the case with $c_{\phi}$ chosen properly. In this case, if ${\tilde{V}}$ in Eq. (2.24) is dominated by the $\lambda_{\phi}\phi^{4}$-term during inflation, in the slow-roll limit of the primordial inflation, the evolution equation in the Jordan frame is approximated as  [^9] Actually, since the dynamics is constrained to be on the ellipse defined in Eq. (2.31), one of the equations in Eq. (2.40) is redundant. Also, in the field-equations there are terms from the field-dependence of $\xi_{\alpha}$. However, those terms are subdominant and have negligible impacts on the field-dynamics. So, we ignored them. [8]
@@
Hence, one can obtain sufficiently large amount of $e$-folds if $X_{2}(0)\gg X_{1}(0)$. Note that, since $X_{1}$ barely changes for a sufficiently small $\xi_{1}$ while $X_{2,3}$ becomes negligible in the end, it is possible to assume that Planck scale is essentially fixed with a value given by
@@
Then, the physics becomes the same as that of nonminimal gravitational couplings of scalar fields other than $\chi$ with the fixed Planck scale. We take this approximation in the subsequent discussion to describe the dynamics $X_{2,3}$ in the more transparent and clear, conventional way.
Once the Planck scale is (nearly) fixed in this way, we can follow the conventional way of converting the Jordan frame to the Einstein frame. The frame-function in Eq. (2.29) is now expressed as
$\Omega=1+\frac{\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\left(1+\alpha\cos(2\theta)\right)=1+\frac{\xi_{r}\phi_{r}^{2}}{M_{\rm P}^{2}}+\frac{\xi_{i}\phi_{i}^{2}}{M_{\rm P}^{2}}\ .$ (2.49)
@@
while the action of the matter sector with scalar fields only is given by
$S_{\rm M}=\int d^{4}x\sqrt{-g}\left\{\frac{1}{2\Omega}\left[\sum_{a}\left(\partial\phi^{a}\right)^{2}\right]-\frac{{\tilde{V}}}{\Omega^{2}}\right\}$ (2.52)
The action for the scalar fields can now be written as
@@
where the symmetric field-space metric $\mathcal{K}_{ab}$ in the basis of $\phi^{a}=\left(\phi_{r},\phi_{i}\right)^{T}$ is given by
$\mathcal{K}_{ab}=\Omega^{-2}\left[\Omega\delta_{ab}+\frac{3M_{\rm P}^{2}}{2}\frac{\partial\Omega}{\partial\phi^{a}}\frac{\partial\Omega}{\partial\phi^{b}}\right]\equiv\Omega^{-2}K_{ab}\ .$ (2.54)
At low energy with $\phi\lll M_{\rm P}$, $\Omega\to 1$ and the kinetic terms return to the canonical form. The shape of the potential $U$ is shown in Fig. 1.
@@
The nonminimal symmetry-breaking gravitational interaction, $\xi_{a}$-term, contributes to the mass of the axi-Majoron, $a_{\phi}\equiv\theta\phi$, the angular mode of $\Phi$ field for a given nonzero $\phi$. If $\xi_{\phi}$ were a constant, in the Einstein frame, the mass-squared of axi-Majoron at low energy obtains a tree-level contribution given by [^10] The low-energy definition of axi-Majoron differs due to the symmetry-breaking nonzero VEVs of the PQ-charged two Higgses, but such a change does not affect our argument.
$\Delta m_{a,\rm tree}^{2}\equiv\frac{\partial^{2}U}{\phi^{2}\partial\theta^{2}}=8\alpha\xi_{\phi}\frac{U}{M_{\rm P}^{2}}\ .$ (2.55)
Although the contribution could be sizable relative to the genuine QCD-contribution, it does not disturb the axion solution, since $U\to 0$ at low energy. However, the symmetry-breaking terms play the key role for baryogenesis after inflation, as discussed in Sec. 4.
Meanwhile, the exchange of massless graviton causes interactions such as $\mathcal{F}\partial^{2}\mathcal{F}$ and $\mathcal{F}\sum\limits_{a}\left(\partial\varphi^{a}\right)^{2}$ [38]. Those interactions result in a loop-induced additional axi-Majoron potential other than the QCD-contribution [^11] We thank to Hyun Min Lee for pointing out this effect.. The axi-Majoron mass from the contribution is
$\Delta m_{a,\rm loop}$ $=$ $\mathcal{O}(10^{-1})\times\alpha^{1/2}\xi_{\phi}\lambda_{\phi}\left(\frac{\phi_{0}^{2}}{M_{\rm P}}\right)\ln^{1/2}\left(\frac{\chi^{2}}{m_{\phi}^{2}}\right)$ (2.56)
$=$ $\mathcal{O}(1)\times\xi_{\phi}\left(\frac{\alpha}{10^{-4}}\right)^{1/2}\left(\frac{\lambda_{\phi}}{6\times 10^{-10}}\right)\left(\frac{\phi_{0}}{10^{12}{\rm GeV}}\right)^{2}\ {\rm keV}\ ,$
where we took $\chi=100M_{\rm P}$ for the numerical estimation in the second line. We will be interested in $\xi_{\phi}\gtrsim\mathcal{O}(1)$ during inflation (see Sec. 3). Eq. (2.56) implies that, if $\xi_{\phi}\gtrsim\mathcal{O}(1)$ even at low energy, the axion solution to the strong CP-problem can work only if $\alpha\ll\mathcal{O}(10^{-20})$ which is too small for baryogenesis in our scenario. This danger is remedy thanks to the field-dependence of $\xi_{\phi}$ as assumed in Eq. (2.7). The coupling is essentially turned off at low energy due to the exponential field-dependence. Hence, the symmetry-breaking nonminimal couplings do not cause the axion-quality problem in our scenario.